a venturimeter is used to measure liquid flow rate of 7.5x103L/min. the difference in pressure across the venturimeter is 8m of the flowing fluid. the pipe diameter is 0.19m. determine the throat diameter of the venturimeter. given Cd value is 0.96


1
Expert's answer
2021-07-12T12:14:00-0400

"Q=C_d\frac{A_1A_2}{\sqrt{A_1^2-A_2^2}}\sqrt{2gh(\frac{\rho_{Hg}}{\rho}-1)},"

"A=\frac{\pi d^2}{4},"

"d_2=\frac4{\pi \sqrt[4]{\frac{2ghC_d^2(\frac{\rho_{Hg}}{\rho}-1)}{Q^2}+\frac{16}{\pi^2 d_1^4}}}=6.9~cm."